Hi, I also don't know if this question goes here but I would like some help.
If f(x)=x^2 and g(x) = x + 2, find f(g(3)) and g(f(3))?
Thank you to anyone who helps me (:

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

If f(x)=x^2 and g(x) = x + 2, find f(g(3)) and g(f(3))?
------------
Find f(3) and g(3)
f(x) = x^2
Sub 3 for x
f(3) = 3^2 = 9
---
g(3) = 3 + 2 = 5
=====================
f(g(3)) = f(5) = 5^2 = 25
g(f(3)) = g(9) = 9 + 2 = 11

Answer by Theo(13342)

f(x) = x^2
g(x) = x+2

f(g(x)) = (x+2)^2
g(f(x)) = x^2+2

f(3) = 3^2 = 9
g(f(3)) = 9+2 = 11

g(3) = 5
f(g(3)) = 5^2 = 25

what you are doing wih the composite function, you are replacing x with f(x)

example 1:

g(x) = x + 2
f(x) = x^2
f(g(x)) = (x+2)^2
what you did was replace the x in f(x) with g(x).

example 2:

f(x) = x^2
g(x) = x + 2
g(f(x)) = x^2 + 2
what you did was replace the x in g(x) with f(x)

here's a reference.

https://www.mathsisfun.com/sets/functions-composition.html
if you have any questions about how this is done, let me know and i'll explain further as best i can.
theo